# Prediction of Energy Resolution in the JUNO Experiment

### v1

@@ 6.2 Event reconstruction methodology and results
The aim of the event reconstruction is to derive the energy and vertex of the event from the charge and time information of photon hits captured by PMTs. However, for large volume LS detectors such as JUNO, various complex optical processes occur during the photon propagation. It is usually rather difficult to build a comprehensive optical model to precisely describe the photon hits of PMTs.
Given that a non-uniform detector response is one of the main contributors to the energy resolution for large LS detectors, precise vertex reconstruction is needed to correct for the energy response non-uniformity. Several algorithms have been developed for vertex reconstruction in JUNO [62, 63]. These algorithms utilize the time information of the first photon hit of PMTs, together with the residual time probability distribution function (pdf), which is mainly determined by the LS timing profile and PMTs’ TTS. Meanwhile, an optical-model-independent method [64] has been developed to reconstruct the event energy in JUNO and was optimized in Ref. [65] to further improve the energy uniformity. The basic principle is to obtain the expected charge for PMTs using calibration data from the automatic calibration unit (ACU) and the cable loop system (CLS), which is then used to build a likelihood function given the observed charge of all the PMTs.
A calibration data driven, simultaneous vertex and energy reconstruction method has been developed for JUNO [66], based on the vertex and energy reconstruction methods described above. The observables used are the charge and time information of the PMTs. Calibration data with known vertex and energy are used to construct the expected charge and time response for each PMT. Given the observed and expected charge and time information of PMTs, a maximum likelihood method is developed to reconstruct the event vertex $\mathbf{r}$ and visible energy $E_{\textrm{}}$ simultaneously, using the likelihood function in Eq. 10.
@@
$&\mu_{i}(\mathbf{r},E_{\textrm{}})=E_{\textrm{}}\times\hat{\mu_{i}}^{L}(r,\theta,\theta_{\text{PMT},i})+\mu_{i}^{D}\\ &t^{j}_{res}=t_{h}^{j}-t^{j}_{tof}(d)-t_{0}\\ \end{aligned}$ (11)
@@
The first product on the right side of Eq. 10 corresponds to the charge-based likelihood function. The index $i$ runs over all the PMTs. The term in the brackets simply describes the probability of observing charge $q_{i}$ on PMT $i$ when the expected number of PE is $\mu_{i}$, and the expected number of PE for the $i$-th PMT which strongly depends on both the vertex and energy of the positron. Given that photons emitted from the same particle have strong temporal correlation and usually arrive on PMTs within a few hundred ns, while PMT dark noise occurs randomly in time, a signal window of 420 ns was set to reduce the PE contamination from dark noise. $\mu_{i}^{D}$ represents the residual PE contribution from dark noise within the signal window. As one of the most crucial ingredients of the reconstruction, $\hat{\mu_{i}}^{L}(r,\theta,\theta_{\textrm{PMT},i})$ represents the expected number of LS PE per unit of visible energy, originating from particles within the signal window, and is obtained using the simulated calibration data in which the calibration source is deployed at various positions in the CD. $P(k,\mu_{i})$ is the Poisson probability for detecting $k$ PE, $P_{Q}(q_{i}|k)$ is the charge pdf of $k$ PE on PMT $i$ that can be constructed by convolving the SPE charge spectrum. Note that for PMTs which do not pass the firing threshold of $q_{i}>0.1\,\textrm{PE}$, this term simplifies to $P(0,\mu_{i})$ =$e^{-\mu_{i}}$. Moreover, the index $k$ ends when $P_{Q}(q_{i}|k)<10^{-8}$ to simplify the calculation.
@@
The second product on the right side of Eq. 10 corresponds to the time-based likelihood function. The index $j$ only runs over the fired PMTs satisfying $-100\,\textrm{ns}<t^{j}_{res}<500\,\textrm{ns}$ and $0.1~{}\textrm{PE}<q_{j}<K$. A cutoff value $K=20$ is set for the detected nPE $k$ to simplify the calculation. The residual hit time $t_{res}$ of PMTs is obtained by subtracting the time of flight $t_{tof}$ and reference time $t_{0}$ from the first hit time $t_{h}$ of PMTs. The distance between the vertex and the PMT is denoted as $d$. $\mu^{l}$ and $\mu^{d}$ stand for the expected number of PE originating from particle or dark noise within the full electronic readout window respectively. Another crucial ingredient of the reconstruction is the residual time PDF $P_{T}(t^{j}_{res}|r,d_{j},\mu^{l}_{j},\mu^{d}_{j},k)$, which is obtained using the same calibration data used for $\hat{\mu_{i}}^{L}$. The fine-grained parameterization of $P_{T}(t^{j}_{res})$ takes into account its dependence on the vertex as well as the distance between the vertex and the PMT. Meanwhile, the impact from dark noise is also included via an analytical approach. More details of the construction of $P_{T}(t^{j}_{res})$ can be found in Ref. [66]. $P(k,\mu^{l}_{j}+\mu^{d}_{j})$ acts as a weight for different $k$ values.
@@
Comparing to previous reconstruction studies, a few important updates should be mentioned. Firstly, two crucial ingredients have been made more realistic: on the one hand, the residual time pdf are derived from calibration data instead of MC simulation, and all PMT electronics effects are considered for the construction of the nPE map. Secondly, the fine-grained parameterization of the residual time pdf as well as the calibration of the time of flight as a function of the photon propagation distance makes the pdf more accurate. Thirdly, the charge and time information of PMTs are combined together to improve the vertex resolution, especially near the acrylic sphere edge. Finally, the vertex and energy are reconstructed simultaneously, which naturally handles the strong correlation between these two quantities.
@@
To evaluate the performance of the energy reconstruction in JUNO, a few sets of positron samples with different kinetic energies $E_{k}$ = (0, 0.5, 1, 2, 3, 4, 5, 6, 8, 11) MeV were produced by MC simulation as summarized in Tab. 3. For each data set, the positrons were uniformly distributed in the CD and the total statistics per set is 500k. In addition, the 68Ge calibration data listed in Tab. 3 were also produced to obtain the expected nPE map and time pdfs of PMTs. The positions and type of calibration source have been slightly optimized based on Ref. [66] to improve the energy uniformity. A realistic detector geometry with all the latest knowledge on the properties of LS and PMTs from previous sections were implemented in the simulation. For each set of positrons with fixed kinetic energy $E^{i}_{k}$, the simultaneous vertex and energy reconstruction using Eq. 10 was applied. The distribution of the reconstructed visible energy $E_{\textrm{rec}}$ was fitted with a Gaussian function ($E^{i}_{\textrm{vis}},\sigma^{i}$) and the results are summarized in Tab. 4. The corresponding energy resolution was defined as the ratio of $\sigma^{i}$/$E^{i}_{\textrm{vis}}$ and the energy non-linearity was calculated by $E^{i}_{\textrm{vis}}$/${E^{i}_{\textrm{dep}}}$ where ${E^{i}_{\textrm{dep}}}$ = ${E^{i}_{k}}+1.022~{}\textrm{MeV}$.
Table 3: List of the MC simulation samples.
@@
68Ge $0.511\cdot 2\,$MeV 20k/point ACU+CLS (293 points)
In Fig. 11, the left plot shows the energy resolution as a function of the average visible energy $E_{\textrm{vis}}$. The data points were fitted with a generic parameterization formula of Eq. 12.
@@
In this equation, $a$ is the statistical term mainly driven by the Poisson statistics of detected PE. The $b$ term is a constant, independent of energy and mostly contributed by the scintillation quenching effect, Cherenkov radiation and energy non-uniformity. The $c$ term accounts for the PMT dark noise and also the positrons’ annihilation $\gamma$s. The best-fit results of $a$, $b$, and $c$ were listed as the Default Case in Tab. 5 and the fitted energy resolution is 2.95% at 1 MeV. The right plot in Fig. 11 shows the energy non-linearity curve. These results were used for the NMO sensitivity in JUNO paper  [14].
@@
Figure 11: Energy resolution (left) and energy non-linearity (right) for positrons using samples from Tab. 3. A fit with Eq. 12 was done for the points in the left plot, while interpolation was used instead in the right plot.
@@
A few additional checks were performed to validate the results. The left plot in Fig. 12 shows that the energy non-uniformity is within 0.4% for positrons with different energies. The right plot shows how the energy resolution changes with respect to $r^{3}$, which is caused mainly by the change of the total number of detected PE.
@@
Figure 12: Energy response non-uniformity checks. Normalized average $E_{\textrm{rec}}$ (left) and energy resolution (right) with respect to $r^{3}$ for positron samples with different energies. The normalization is done by dividing the average value within the FV. The Red vertical line corresponds to the FV cut.
7 Decomposition of the energy resolution
@@
The contribution of major factors to the energy resolution budget is estimated at two different phases: the detector simulation phase and the calibration and reconstruction process. In the detector simulation phase, positron samples are generated at the CD center. However, in the reconstruction process, the positron vertexes are generated with a uniform distribution within the LS volume.
7.1 Decomposition of the energy resolution in the detector simulation
In the updated detector simulation, the predicted PE yield of the JUNO detector is 1665 PE/MeV at the detector center, which is larger compared to the previously reported value of 1345 PE/MeV [24]. This enhancement is attributed to several factors, mainly higher LPMTs’ PDE ($\sim$11%), new CD geometry with a reflective optical mask on PMT modules ($\sim$3%), and a new PMT optical model ($\sim$8%) discussed in Sec. 4.6. Furthermore, the Cherenkov to total PE ratio is found to be 1.1% for 1 MeV positrons, and the detector simulation indicates a PE yield of 1785 PE/MeV for uniformly distributed events of neutron capture on hydrogen.
@@ 7.1 Decomposition of the energy resolution in the detector simulation
The positron energy resolution at the center of the CD is obtained from detector simulation. According to the recorded truth information, the contributions from scintillation light, Cherenkov light, and their covariance to the energy resolution are extracted and major effects in the energy resolution budget are shown in Fig. 13. It has been previously discussed that the quenching effect causes the number of scintillation PE to deviate from a Poisson distribution. Therefore, the contribution from scintillation PE can be further decomposed into two parts: the Poisson fluctuation (blue curve) and the quenching effect (red curve). The contribution from Poisson fluctuation is evaluated using the square root of the number of scintillation PE, which is given by $\sigma_{\textrm{stat}}=\sqrt{N_{\textrm{ScintPE}}}$. This component represents the statistical fluctuations in the scintillation process. The contribution from the quenching effect is then obtained by subtracting the Poisson standard deviation from the standard deviation of the scintillation PE number distribution. Mathematically, this is expressed as $\sigma_{\textrm{quench}}=\sqrt{\sigma_{\textrm{ScintPE}}^{2}-\sigma_{\textrm{stat}}^{2}}$, where $\sigma_{\textrm{ScintPE}}$ is the standard deviation of the scintillation PE number distribution. In this figure, the yellow curve represents the contribution from Cherenkov radiation, and the dark red curve represents the correlation between the scintillation and Cherenkov processes. As expected, the gray curve, which is the square root of the quadratic sum of the four components, accurately reproduces the total energy resolution (black curve) obtained directly from the standard deviation of the total PE spectrum.
@@
Figure 13: Energy resolution decomposition. The black curve refers to the energy resolution obtained with total PE. The blue, red, dark red, and yellow curves represent the contributions from scintillation Poisson statistics, quenching effect, Cherenkov radiation and the covariance between scintillation, and Cherenkov processes, respectively. The gray curve which superimposes with the black curve is the root squared sum of the four components.
@@
7.2 Decomposition of the energy resolution in the reconstruction
There are a few key factors which could affect the energy resolution in the event reconstruction. Due to energy non-uniformity, the vertex uncertainty will propagate to the energy resolution. A more precise vertex could improve the energy resolution. Another important contributor comes from the PMT dark noise, which contaminates the signal photons and worsens the energy resolution. Thirdly, the inputs to the event reconstruction, namely the charge and time of PMTs which are obtained from the PMT waveform reconstruction, the charge non-linearity and waveform reconstruction uncertainty will propagate to the energy resolution as well. Finally, the PMT charge is used to estimate the number of detected PE. The intrinsic PMT charge resolution, which includes contributions from both Gaussian and exponential components (Sec. 5.1.1), will lead to charge smearing and potentially worsen the energy resolution for the charge-based energy reconstruction. The impact of these factors on the energy resolution is decomposed in this section. Different cases listed below were compared, in which the factors above were removed sequentially.
@@ 7.2 Decomposition of the energy resolution in the reconstruction
  • Case C: on top of Case B, the waveform reconstruction is removed in the reconstruction.
@@
In the default case, the positron vertex is simultaneously reconstructed with its energy using the method described in Sec. 6.2. The vertex resolution is around 10 cm at 1 MeV and decreases at higher energies. Meanwhile, the vertex bias is less than 2 cm within the FV. In case A everything is the same as the default case, except that the MC truth vertex is used in the energy reconstruction, which is equivalent to a vertex resolution of 0 mm. Besides using the true vertex, the PMT dark noise is removed in case B. Given that there is no more dark noise, the 420 ns signal window is discarded in both the construction of the nPE maps as well as in the energy reconstruction. In case C, the waveform reconstruction is replaced by a toy simulation, which smears the charge of each PE, provides charge and time. This smeared PMT charge from the toy simulation is used in both the construction of the nPE maps as well as the energy reconstruction. In comparison, the charge reconstructed from PMT waveforms are used in previous cases. By using these toy electronic simulation samples, we are able to remove the impact from the waveform reconstruction as well as other electronics effects except the PMT SPE charge smearing. The true vertex is used for case C. For case D, the PMT SPE charge smearing is removed on top of case C by directly using the detector simulation samples for the reconstruction. As a consequence, the new observable is the number of PE, instead of the reconstructed charge. This corresponds to the most ideal scenario and leads to the best energy resolution.
@@
Figure 14: Comparison of the energy resolution (left) and energy non-linearity (right) among the different cases. The energy resolution for each case is better with respect to the previous one, indicating that the energy resolution improves by moving from the real to the ideal situation.
@@
Figure 14 shows the comparison of the energy resolution and energy non-linearity among all the cases. We can see that the energy resolution improves monotonically among the cases, indicating that the energy resolution keeps improving by removing each contributing factor. The fitting results of the energy resolution for all the cases are summarized in Tab. 5. The relative improvement of the energy resolution at 1 MeV for each case with respect to the previous case is also presented. We can see that at 1 MeV, the PMT dark noise and SPE charge smearing are the two dominant contributing factors to the energy resolution with a relative improvement of about 3.45% and 2.1%, respectively. Removing the charge uncertainty from the PMT waveform reconstruction together with charge non-linearity leads to a 0.92% relative improvement. The impact of vertex resolution is relatively small, the relative improvement is about  0.78% using the true vertex.
@@
Table 5: Comparison of the energy resolution among all the cases. The relative improvement of the energy resolution at 1 MeV with respect to the previous case is also shown in the last column.
@@
The comparison above shows the average impact of each factor on the energy resolution. In addition, it also enables us to decompose the energy resolution at each discrete energy point. With the energy resolution curves of all the cases, the default energy resolution can be decomposed into five components, as it is shown in Fig. 15. The black solid curve corresponds to the default case, which is used for the NMO analysis [14]. The teal solid curve corresponds to case D and represents the ideal energy resolution we could obtain. The four dashed curves show the contribution of each factor to the energy resolution, and the values at each energy point are calculated as $\sqrt{(E^{X}_{\textrm{res}})^{2}-(E^{Y}_{\textrm{res}})^{2}}$, where Y goes from case A to D and X corresponds to the preceding case Y. Two features stand out immediately: first, the contribution from dark noise decreases as the energy increases as expected. Moreover, in the energy range below 1.5 MeV, dark noise is the major contributor. For the energy range above 1.5 MeV, particularly the most sensitive region (1.5, 3) MeV for the NMO analysis, the SPE charge smearing is the dominant factor. Preliminary studies indicate that its impact could be partially mitigated by PE counting, especially for PMTs with 1 or 2 PE.
Figure 15: Decomposition of the energy resolution. The black solid curve represents the default case. It is decomposed into five parts: the four dotted curves correspond to the contribution of vertex resolution (red), dark noise (green), waveform reconstruction (blue), and SPE charge smear (purple) respectively. The light blue solid curve represents case D (ideal case). Taking 1.022 MeV as an example, the values for all the curves are 3.12%, 0.35%, 0.83%, 0.39%, 0.58% and 2.90%, and the decomposition can be written as $3.12^{2}=0.35^{2}+0.83^{2}+0.39^{2}+0.58^{2}+2.90^{2}$.

### v2

@@ 6.2 Event reconstruction methodology and results
The aim of the event reconstruction is to derive the energy and vertex of the event from the charge and time information of photon hits captured by PMTs. However, for large-volume LS detectors such as JUNO, various complex optical processes occur during the photon propagation. It is usually relatively difficult to build a comprehensive optical model to precisely describe the photon hits of PMTs.
Given that a non-uniform detector response is one of the main contributors to the energy resolution for large LS detectors, precise vertex reconstruction is needed to correct for the energy response non-uniformity. Several algorithms have been developed for vertex reconstruction in JUNO [69, 70]. These algorithms utilize the time information of the first photon hit of PMTs, together with the residual time probability distribution function (pdf), which is mainly determined by the LS timing profile and PMTs’ TTS. Meanwhile, an optical-model-independent method [71] was developed to reconstruct the event energy in JUNO and was optimized in Ref. [72] to further improve the energy uniformity. The basic principle is to obtain the expected charge for PMTs using calibration data from the automatic calibration unit (ACU) and cable loop system (CLS), which is then used to build a likelihood function given the observed charge of all PMTs.
A calibration data-driven simultaneous vertex and energy reconstruction method has been developed for JUNO [73], based on the vertex and energy reconstruction methods described above. The observables considered are the charge and time information of the PMTs. Calibration data with known vertices and energy are used to construct the expected charge and time response for each PMT. Given the observed and expected charge and time information of PMTs, a maximum likelihood method is developed to reconstruct the event vertex $\mathbf{r}$=$\vec{r}(r,\theta,\phi)$ and visible energy $E_{\textrm{}}$ simultaneously, using the likelihood function in Eq. 10:
@@
$&\mu_{i}(\mathbf{r},E_{\textrm{}})=E_{\textrm{}}\times\hat{\mu_{i}}^{L}(r,\theta,\theta_{\text{PMT},i})+\mu_{i}^{D}\\ &t^{j}_{res}=t^{j}-t^{j}_{tof}(d_{j})-t_{0}\\ \end{aligned}$ (11)
@@
The first product on the right side of Eq. 10 corresponds to the charge-based likelihood function. The index $i$ runs over all PMTs. The term in parentheses simply describes the probability of observing charge $q_{i}$ on PMT $i$ when the expected number of PEs is $\mu_{i}$, which strongly depends on both the vertex and energy of the positron. Given that photons emitted from the same particle have strong temporal correlation and usually arrive on PMTs within a few hundred ns, while PMT dark noise occurs randomly in time, a signal window of 420 ns is set to reduce the PE contamination from dark noise. $\mu_{i}^{D}$ represents the residual PE contribution from dark noise within the signal window. As one of the most crucial ingredients of the reconstruction, $\hat{\mu_{i}}^{L}(r,\theta,\theta_{\textrm{PMT},i})$ represents the expected number of LS PEs per unit of visible energy, originating from particles within the signal window. This is obtained using the simulated calibration data, in which the calibration source is deployed at various positions in the CD. Here, $r$ and $\theta$ are the components of the particle vertex $\mathbf{r}$, while $\theta_{\textrm{PMT},i}$ is the angle between $\mathbf{r}$ and the PMT position vector $\mathbf{r}_{\textrm{PMT},i}$. $P(k,\mu_{i})$ is the Poisson probability for detecting $k$ PEs, $P_{Q}(q_{i}|k)$ is the probability of observing charge $q_{i}$ on PMT $i$ given the charge pdf of $k$ PE $P_{Q}(q|k)$, which can be constructed by convolving the SPE charge spectrum with $P_{Q}(q|k-1)$. Note that for PMTs that do not pass the firing threshold of $q_{i}>0.1~{}\textrm{PE}$, this term simplifies to $P(0,\mu_{i})$+$P_{Q}(q_{i}<0.1PE|k=1)$*$P(1,\mu_{i})$. Moreover, the index $k$ ends when $P_{Q}(q_{i}|k)<10^{-8}$ is met to simplify the calculation.
@@
The second product on the right side of Eq. 10 corresponds to the time-based likelihood function. The index $j$ only runs over the fired PMTs satisfying $-100~{}\textrm{ns}<t^{j}_{res}<500~{}\textrm{ns}$ and $0.1~{}\textrm{PE}<q_{j}<K$. A cutoff value of $K=20$ is set for the detected nPE $k$ to simplify the calculation. The residual hit time $t_{res}$ of PMTs is obtained by subtracting the time of flight $t_{tof}$ and reference time $t_{0}$ from the first hit time $t$ of PMTs. The distance between the vertex and PMT is denoted as $d$. $\mu^{l}$ and $\mu^{d}$ represent the expected number of PEs originating from particle or dark noise within the full electronic readout window, respectively. Another crucial ingredient of the reconstruction is the residual time PDF $P_{T}(t^{j}_{res}|r,d_{j},\mu^{l}_{j},\mu^{d}_{j},k)$, which is obtained using the same calibration data used for $\hat{\mu_{i}}^{L}$. The fine-grained parameterization of $P_{T}(t^{j}_{res})$ takes into account its dependence on the vertex radius as well as the distance between the vertex and PMT. Meanwhile, the impact from dark noise is also included via an analytical approach. More details of the construction of $P_{T}(t^{j}_{res})$ can be found in Ref. [73]. $P(k,\mu^{l}_{j}+\mu^{d}_{j})$ acts as a weight for different $k$ values.
@@
Compared to previous reconstruction studies, a few important updates should be mentioned. First, two crucial ingredients have been made more realistic: the residual time pdfs are derived from calibration data instead of MC simulation, and all PMT electronics effects are considered for the construction of the expected nPE map $\mu_{i}$. Second, the fine-grained parameterization of the residual time pdf as well as the calibration of the time of flight as a function of the photon propagation distance makes the pdf more accurate. Third, the charge and time information of PMTs are combined together to improve the vertex resolution, especially near the acrylic sphere edge. Finally, the vertex and energy are reconstructed simultaneously, which naturally handles the strong correlation between these two quantities.
@@
To evaluate the performance of the energy reconstruction in JUNO, a few sets of positron samples with different kinetic energies $E_{k}$ = (0, 0.5, 1, 2, 3, 4, 5, 6, 8, 11) MeV were produced by MC simulation, as summarized in Tab. 3. For each data-set, the positrons are uniformly distributed in the CD, and the total statistics per set is 500k. In addition, the 68Ge calibration data listed in Tab. 3 were also produced to obtain the expected nPE map and time pdfs of PMTs. The positions and type of calibration source have been slightly optimized based on Ref. [72] to improve the energy uniformity. A realistic detector geometry with all the latest knowledge on the properties of LS and PMTs from previous sections were implemented in the simulation. For each set of positrons with fixed kinetic energy $E^{i}_{k}$, the simultaneous vertex and energy reconstruction using Eq. 10 was applied. The distribution of the reconstructed visible energy $E_{\textrm{rec}}$ was fitted with a Gaussian function ($E^{i}_{\textrm{vis}},\sigma^{i}$), and the results are summarized in Tab. 4. The corresponding energy resolution is defined as the ratio of $\sigma^{i}$/$E^{i}_{\textrm{vis}}$, and the energy non-linearity is calculated by $E^{i}_{\textrm{vis}}$/${E^{i}_{\textrm{dep}}}$, where ${E^{i}_{\textrm{dep}}}$ = ${E^{i}_{k}}+1.022~{}\textrm{MeV}$.
Table 3: List of MC simulation samples.
@@
68Ge $0.511\cdot 2~{}$MeV 20k/point ACU+CLS (293 points)
In Fig. 11, the left plot shows the energy resolution as a function of the average visible energy $E_{\textrm{vis}}$. The data points are fitted with a generic parameterization formula as follows:
@@
In this equation, $a$ is the statistical term mainly driven by the Poisson statistics of detected PE. The $b$ term is a constant, independent of energy and mostly contributed by the scintillation quenching effect, Cherenkov radiation, and energy non-uniformity. The $c$ term accounts for the PMT dark noise and the positron annihilation $\gamma$s. The best-fit results of $a$, $b$, and $c$ are listed as the Default Case in Tab. 5, and the fitted energy resolution is 2.95% at 1 MeV. The right plot in Fig. 11 shows the energy non-linearity curve. These results were used for the NMO sensitivity in the JUNO paper  [18].
@@
Figure 11: Energy resolution (left) and energy non-linearity (right) for positrons using samples from Tab. 3. A fit with Eq. 12 was performed for the points in the left plot, while interpolation was used instead in the right plot.
@@
A few additional checks were performed to validate the results. The left plot in Fig. 12 shows that the energy non-uniformity is within 0.4% for positrons with different energies. The right plot shows how the energy resolution changes with respect to $r^{3}$, which is caused mainly by the change in the total number of detected PEs.
@@
Figure 12: Energy response non-uniformity checks. Normalized average $E_{\textrm{rec}}$ (left) and energy resolution (right) with respect to $r^{3}$ for positron samples with different energies. The normalization is conducted by dividing the average value within the FV. The red vertical line corresponds to the FV cut.
7 Decomposition of energy resolution
@@
The contribution of major factors to the energy resolution budget is estimated at two different phases: detector simulation phase and calibration and reconstruction process. In the detector simulation phase, positron samples are generated at the CD center. However, in the reconstruction process, the positron vertexes are generated with a uniform distribution within the LS volume.
7.1 Decomposition of energy resolution in the detector simulation
@@
The positron energy resolution at the center of the CD is obtained from detector simulation. According to the recorded truth information, the contributions from scintillation light, Cherenkov light, and their covariance to the energy resolution are extracted, and major effects in the energy resolution budget are shown in Fig. 13. It has previously been discussed that the quenching effect causes the number of scintillation PEs to deviate from a Poisson distribution. Therefore, the contribution from scintillation PEs can be further decomposed into two parts: the Poisson fluctuation (blue curve) and quenching effect (red curve). The contribution from Poisson fluctuation is evaluated using the square root of the number of scintillation PEs, which is given by $\sigma_{\textrm{stat}}=\sqrt{N_{\textrm{ScintPE}}}$. This component represents the statistical fluctuations in the scintillation process. The contribution from the quenching effect is then obtained by subtracting the Poisson standard deviation from the standard deviation of the scintillation PE number distribution. Mathematically, this is expressed as $\sigma_{\textrm{quench}}=\sqrt{\sigma_{\textrm{ScintPE}}^{2}-\sigma_{\textrm{stat}}^{2}}$, where $\sigma_{\textrm{ScintPE}}$ is the standard deviation of the scintillation PE number distribution. In Fig. 13, the yellow curve represents the contribution from Cherenkov radiation, and the dark red curve represents the correlation between the scintillation and Cherenkov processes. As expected, the gray curve, which is the square root of the quadratic sum of the four components, accurately reproduces the total energy resolution (black curve) obtained directly from the standard deviation of the total PE spectrum.
@@
Figure 13: Energy resolution decomposition. The black curve represents the energy resolution obtained with total PE. The blue, red, dark red, and yellow curves represent the contributions from scintillation Poisson statistics, quenching effect, Cherenkov radiation and the covariance between scintillation, and Cherenkov processes, respectively. The gray curve, which superimposes with the black curve, is the root squared sum of the four components.
@@
7.2 Decomposition of energy resolution in the reconstruction
There are a few key factors that could affect the energy resolution in the event reconstruction. Due to energy non-uniformity, the vertex uncertainty will propagate to the energy resolution. A more precise vertex could improve the energy resolution. Another important contributor comes from the PMT dark noise, which contaminates the signal photons and worsens the energy resolution. Moreover, the inputs to the event reconstruction, namely, the charge and time of PMTs, are obtained from the PMT waveform reconstruction, charge non-linearity and waveform reconstruction uncertainty, will propagate to the energy resolution as well. Finally, the PMT charge could only provide a rough estimation of the number of detected PEs. The intrinsic PMT charge resolution, which includes contributions from both Gaussian and exponential components (Sec. 5.1.1), will lead to charge smearing and potentially worsen the energy resolution for the charge-based energy reconstruction. The impact of these factors on the energy resolution is decomposed in this section. Different cases were considered, in which the factors above were removed sequentially, as follows:
@@ 7.2 Decomposition of energy resolution in the reconstruction
  • Case C: on top of Case B, the waveform reconstruction is replaced by a toy simulation to provide the PMT charge and time in the reconstruction.
@@
In the default case, the positron vertex is simultaneously reconstructed with its energy using the method described in Sec. 6.2. The vertex resolution is approximately 10 cm at 1 MeV and decreases at higher energies. Meanwhile, the vertex bias is less than 2 cm within the FV. In case A, everything is the same as in the default case, except that the MC truth vertex is used in the energy reconstruction, which is equivalent to a vertex resolution of 0 mm. Besides using the true vertex, the PMT dark noise is removed in case B. Given that there is no more dark noise, the 420 ns signal window is discarded in both the nPE map and energy reconstruction. In case C, the waveform reconstruction is replaced by a toy simulation that smears the charge of each PE, providing both charge and time information. This smeared PMT charge from the toy simulation is used in both the nPE map and the energy reconstruction. In comparison, the charge reconstructed from PMT waveforms is used in previous cases. By using these toy electronic simulation samples, we are able to remove the impact from the waveform reconstruction as well as other electronics effects, except the PMT SPE charge smearing. The true vertex is used for case C. For case D, the PMT SPE charge smearing is removed on top of case C by directly using the detector simulation samples for the reconstruction. Consequently, the new observable is the PE number, instead of the reconstructed charge. This corresponds to the most ideal scenario and leads to the best energy resolution.
@@
Figure 14: Comparison of the energy resolution (left) and energy non-linearity (right) among the different cases. The energy resolution for each case is better than that of the previous one, indicating that the energy resolution improves by moving from the real to ideal situation. The black curve from Fig. 13 is represented by the dashed curve here, and its corresponding energy resolution is close to that of case D.
@@
Figure 14 shows a comparison of the energy resolution and energy non-linearity among all cases. We can see that the energy resolution improves monotonically among the cases, indicating that the energy resolution keeps improving by removing each contributing factor. The fitting results of the energy resolution for all cases are summarized in Tab. 5. The relative improvement of the energy resolution at 1 MeV for each case with respect to the previous case is also presented. We can see that at 1 MeV, the PMT dark noise and SPE charge smearing are the two dominant contributing factors to the energy resolution, with relative improvements of approximately 3.45% and 2.1%, respectively. Removing the charge uncertainty from the PMT waveform reconstruction together with charge non-linearity leads to a 0.92% relative improvement. The impact of vertex resolution is relatively small, and the relative improvement is approximately  0.78% using the true vertex. The black curve from Fig. 13 is also shown as the dashed curve here, and its corresponding energy resolution is close to that of case D.
@@
Table 5: Comparison of the energy resolutions among all the cases. Here a, b and c correspond to the three parameters from Eq. 12. The relative improvement of the energy resolution at 1 MeV with respect to the previous case is also shown in the last column.
@@
The above comparison shows the average impact of each factor on the energy resolution. In addition, it also enables us to decompose the energy resolution at each discrete energy point. With the energy resolution curves of all cases, the default energy resolution can be decomposed into five components, as shown in Fig. 15. The black solid curve corresponds to the default case, which is used for the NMO analysis [18]. The teal solid curve corresponds to case D and represents the ideal energy resolution we could obtain. The four dashed curves show the contribution of each factor to the energy resolution, and the values at each energy point are calculated as $\sqrt{(E^{X}_{\textrm{res}})^{2}-(E^{Y}_{\textrm{res}})^{2}}$, where Y goes from case A to D and X corresponds to the preceding case. Two features stand out immediately. The contribution from dark noise decreases as the energy increases, as expected. Moreover, in the energy range below 1.5 MeV, dark noise is the major contributor. For the energy range above 1.5 MeV, particularly the most sensitive region of (1.5, 3) MeV for the NMO analysis, the SPE charge smearing is the dominant factor. Preliminary studies indicate that its impact could be partially mitigated by PE counting, especially for PMTs with 1 or 2 PEs.
Figure 15: Decomposition of the energy resolution. The black solid curve represents the default case. It is decomposed into five parts: the four dotted curves correspond to the contributions of vertex resolution (red), dark noise (green), waveform reconstruction (blue), and SPE charge smear (purple), and the light blue solid curve represents case D (ideal case). Taking 1.022 MeV as an example, the values for all curves are 3.12%, 0.35%, 0.83%, 0.39%, 0.58% and 2.90%, and the decomposition can be written as $3.12^{2}=0.35^{2}+0.83^{2}+0.39^{2}+0.58^{2}+2.90^{2}$.